# Aethorix v1.0: An Integrated Scientific AI Agent for Scalable Inorganic Materials Innovation and Industrial Implementation

### v2

@@ 9.3 Adaptive Control of Process Parameters
The data scarcity typical of industrial contexts can yield an ill-posed inverse problem for model training. In our example, the trained empirical models exhibited reasonable accuracy for forecasting 1-day compressive strength, but showed significant root mean squared errors (RMSEs) for 3-day and 28-day results. In contrast, the multimodal models reduced the RMSEs by 20-30% for 3-day and 28-day models, elevating all conservative tolerance threshold ($\pm$ 1.5 MPa) accuracy to above 85% and stringent precision threshold ($\pm$ 1.0 MPa) accuracy to above 75%. The multimodel methods incorporated graph representations of predicted dominant clinker and post-hydrated gel phases, as well as scalar value of the corresponding formation energy at 0 K for each phase. The elemental reference energy was shifted in accordance to the stoichiometric ratios of input oxides, as given by:
@@
where $x_{el}$ and $x_{Ca}$ denote the molar compositions of element $x$ and Ca, respectively. Upon completing model training, Bayesian optimization was employed to find process parameters that maximize cement compressive strength across all curing periods (1-day, 3-day, and 28-day) by iteratively refining significant parameters within physically realistic bounds. The optimized parameters identified for maximizing compressive strength are shown in Figure 5d, which quantifies the percentage deviation of these parameter values relative to the baseline operational averages employed in the past production. Strategic production decisions for future production can be derived from these optimization results. As examples, systematic reduction of Al content demonstrates potential for quality enhancement across all curing periods, while S content requires careful balancing. Elevated concentrations benefit 3-day strength development but inversely impact 28-day performance. It is worth noting that Figure 5d shows that the maximum strength would be achieved with lowered K content across all curing periods, contrary to atomistic predictions of phase stabilization through K-doping. The discrepancy may originate from limitations in the MLIP’s accuracy for dopant modeling (see Figure4d) or from kinetic barriers, such as inhibited K transport during clinker formation, which prevent the realization of thermodynamically stable phases. While this necessitates iterative refinement of our proprietary predictive models, a detailed discussion of further development falls beyond the scope of this paper.

### v3

@@ 9.3 Adaptive Control of Process Parameters
The high-quality data scarcity typical of industrial contexts can yield an ill-posed inverse problem for model training. In our example, the empirical models trained on the small dataset exhibited reasonable accuracy for forecasting 1-day compressive strength, but showed significant root mean squared errors (RMSEs) for 3-day and 28-day results. In contrast, the multimodal models reduced the RMSEs by 20-30% for 3-day and 28-day models, elevating all conservative tolerance threshold ($\pm$ 1.5 MPa) accuracy to above 85% and stringent precision threshold ($\pm$ 1.0 MPa) accuracy to above 75%. The multimodel methods incorporated graph representations of predicted dominant clinker and post-hydrated gel phases, as well as scalar value of the corresponding formation energy at 0 K for each phase. The elemental reference energy was shifted in accordance to the stoichiometric ratios of input oxides, as given by:
@@
where $x_{el}$ and $x_{Ca}$ denote the molar compositions of element $x$ and Ca, respectively. The formalism rests on the assumption of ideal solution behavior. Upon completing model training, Bayesian optimization was employed to find process parameters that maximize cement compressive strength across all curing periods (1-day, 3-day, and 28-day) by iteratively refining significant parameters within physically realistic bounds. The optimized parameters identified for maximizing compressive strength are shown in Figure 5d, which quantifies the percentage deviation of these parameter values relative to the baseline operational averages employed in the past production. Strategic production decisions for future production can be derived from these optimization results. As examples, systematic reduction of Al content demonstrates potential for quality enhancement across all curing periods, while S content requires careful balancing. Elevated concentrations benefit 3-day strength development but inversely impact 28-day performance. It is worth noting that Figure 5d shows that the maximum strength would be achieved with lowered K content across all curing periods, contrary to atomistic predictions of phase stabilization through K-doping. The discrepancy may originate from limitations in the MLIP’s accuracy for dopant modeling (see Figure4d) or from kinetic barriers, such as inhibited K transport during clinker formation, which prevent the realization of thermodynamically stable phases. While this necessitates iterative refinement of our proprietary predictive models, a detailed discussion of further development falls beyond the scope of this paper.